MODELING N_{2} AND H_{2} BROADENED CH_{4} INFRARED ABSORPTION SPECTRA. LINE-MIXING MODEL AND COMPARISONS WITH MEASUREMENTS

\underline{HA TRAN} and JEAN-MICHEL HARTMANN, LISA-Faculté des Sciences et Technologie, 94010 Créteil Cedex, France.

Line-mixing effetcs have been studied in ν_{3} and ν_{4} bands of CH_{4} perturbed by N_{2} and H_{2} at various temperatures. The absorption in the 1100-1500 and 2800-3200 cm^{−1} spectral regions for total pressures in the 0.25-2 and 25-500 atm ranges have been measured. A line-mixing model is proposed wich is based on state-to-state collisional transfert rates calculated from intermolecular potential surface with a semiclassical approach. The line-coupling relaxation matrix is constructed from these state-to-state collisional transfert rates and some empirical parameters. The results of comparisons between measurements and spectra computed accounting for and neglecting line-mixing prove the satisfactory quality of the model. Calculations with the line-mixing model of atmospheric spectra are also made.